# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 9. Acknowledgments
- Niu et al. (2024) Zhenxing Niu, Haodong Ren, Xinbo Gao, Gang Hua, and Rong Jin. 2024. Jailbreaking attack against multimodal large language model. arXiv preprint arXiv:2402.02309 (2024).
- NousResearch (2023) NousResearch. 2023. Nous Hermes 2 - Yi-34B. (2023). https://huggingface.co/NousResearch/Nous-Hermes-2-Yi-34B
- OpenAI (2022) OpenAI. 2022. Moderation API. (2022). https://platform.openai.com/docs/guides/moderation/overview
- OpenAI (2023a) OpenAI. 2023a. GPT-4o API. (2023). https://platform.openai.com/docs/models/gpt-4o
- OpenAI (2023b) OpenAI. 2023b. Gpt-4v(ision) system card. (2023). https://cdn.openai.com/papers/GPTV_System_Card.pdf
- OpenAI (2023c) OpenAI. 2023c. Using GPT-4 for content moderation. (2023). https://openai.com/index/using-gpt-4-for-content-moderation/
- OpenAI (2024) OpenAI. 2024. OpenAI Usage policies. (2024). https://openai.com/policies/usage-policies
- OpenBMB (2024) OpenBMB. 2024. OmniLMM. (2024). https://github.com/OpenBMB/MiniCPM-V/blob/main/omnilmm_en.md
@@
- Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. NIPS (2022).
- Papernot et al. (2016a) Nicolas Papernot, Patrick McDaniel, and Ian Goodfellow. 2016a. Transferability in machine learning: from phenomena to black-box attacks using adversarial samples. arXiv preprint arXiv:1605.07277 (2016).
- Papernot et al. (2017) Nicolas Papernot, Patrick McDaniel, Ian Goodfellow, Somesh Jha, Z Berkay Celik, and Ananthram Swami. 2017. Practical black-box attacks against machine learning. In ACM ASIACCS.
- Papernot et al. (2016b) Nicolas Papernot, Patrick McDaniel, Somesh Jha, Matt Fredrikson, Z Berkay Celik, and Ananthram Swami. 2016b. The limitations of deep learning in adversarial settings. In EuroS&P.
@@
- Papernot et al. (2016c) Nicolas Papernot, Patrick McDaniel, Arunesh Sinha, and Michael Wellman. 2016c. Towards the science of security and privacy in machine learning. arXiv preprint arXiv:1611.03814 (2016).
@@
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th annual meeting of the Association for Computational Linguistics.
@@
- Peng et al. (2018) Yuxin Peng, Jinwei Qi, and Yuxin Yuan. 2018. Modality-specific cross-modal similarity measurement with recurrent attention network. TIP (2018).
- Perez and Ribeiro (2022) Fábio Perez and Ian Ribeiro. 2022. Ignore previous prompt: Attack techniques for language models. arXiv preprint arXiv:2211.09527 (2022).
- Phute et al. (2023) Mansi Phute, Alec Helbling, Matthew Hull, ShengYun Peng, Sebastian Szyller, Cory Cornelius, and Duen Horng Chau. 2023. Llm self defense: By self examination, llms know they are being tricked. arXiv preprint arXiv:2308.07308 (2023).
- Pintor et al. (2021) Maura Pintor, Fabio Roli, Wieland Brendel, and Battista Biggio. 2021. Fast minimum-norm adversarial attacks through adaptive norm constraints. NIPS (2021).
- Plummer et al. (2015) Bryan A Plummer, Liwei Wang, Chris M Cervantes, Juan C Caicedo, Julia Hockenmaier, and Svetlana Lazebnik. 2015. Flickr30k entities: Collecting region-to-phrase correspondences for richer image-to-sentence models. In ICCV.
@@
- Qi et al. (2024) Xiangyu Qi, Kaixuan Huang, Ashwinee Panda, Peter Henderson, Mengdi Wang, and Prateek Mittal. 2024. Visual adversarial examples jailbreak aligned large language models. In AAAI.
- Qi et al. (2023) Xiangyu Qi, Yi Zeng, Tinghao Xie, Pin-Yu Chen, Ruoxi Jia, Prateek Mittal, and Peter Henderson. 2023. Fine-tuning aligned language models compromises safety, even when users do not intend to! arXiv preprint arXiv:2310.03693 (2023).
- Qin et al. (2022) Zeyu Qin, Yanbo Fan, Yi Liu, Li Shen, Yong Zhang, Jue Wang, and Baoyuan Wu. 2022. Boosting the transferability of adversarial attacks with reverse adversarial perturbation. NIPS (2022).
- Qraitem et al. (2024) Maan Qraitem, Nazia Tasnim, Kate Saenko, and Bryan A Plummer. 2024. Vision-llms can fool themselves with self-generated typographic attacks. arXiv preprint arXiv:2402.00626 (2024).
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al. 2021. Learning transferable visual models from natural language supervision. In ICML.
- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog (2019).
- Rashtchian et al. (2010) Cyrus Rashtchian, Peter Young, Micah Hodosh, and Julia Hockenmaier. 2010. Collecting image annotations using amazon’s mechanical turk. In NAACL-HLT Workshop.
- Rasiwasia et al. (2010) Nikhil Rasiwasia, Jose Costa Pereira, Emanuele Coviello, Gabriel Doyle, Gert RG Lanckriet, Roger Levy, and Nuno Vasconcelos. 2010. A new approach to cross-modal multimedia retrieval. In ACM MM.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-resolution image synthesis with latent diffusion models. In CVPR.
- Rosen (1960) Jo Bo Rosen. 1960. The gradient projection method for nonlinear programming. Part I. Linear constraints. Journal of the society for industrial and applied mathematics (1960).
- Röttger et al. (2023) Paul Röttger, Hannah Rose Kirk, Bertie Vidgen, Giuseppe Attanasio, Federico Bianchi, and Dirk Hovy. 2023. Xstest: A test suite for identifying exaggerated safety behaviours in large language models. arXiv preprint arXiv:2308.01263 (2023).
- Rudin et al. (1992) Leonid I Rudin, Stanley Osher, and Emad Fatemi. 1992. Nonlinear total variation based noise removal algorithms. Physica D: nonlinear phenomena (1992).
- Russakovsky et al. (2015) Olga Russakovsky, Jia Deng, Hao Su, Jonathan Krause, Sanjeev Satheesh, Sean Ma, Zhiheng Huang, Andrej Karpathy, Aditya Khosla, Michael Bernstein, et al. 2015. Imagenet large scale visual recognition challenge. IJCV (2015).
- Salman et al. (2019) Hadi Salman, Jerry Li, Ilya Razenshteyn, Pengchuan Zhang, Huan Zhang, Sebastien Bubeck, and Greg Yang. 2019. Provably robust deep learning via adversarially trained smoothed classifiers. NIPS (2019).
@@
- SessionGloomy (2023) SessionGloomy. 2023. New jailbreak! Proudly unveiling the tried and tested DAN 5.0. (2023). https://www.reddit.com/r/ChatGPT/comments/10tevu1/new_jailbreak_proudly_unveiling_the_tried_and/
- Shamsabadi et al. (2021) Ali Shahin Shamsabadi, Changjae Oh, and Andrea Cavallaro. 2021. Semantically adversarial learnable filters. TIP (2021).
@@
- Shan et al. (2023) Shawn Shan, Jenna Cryan, Emily Wenger, Haitao Zheng, Rana Hanocka, and Ben Y Zhao. 2023. Glaze: Protecting artists from style mimicry by Text-to-Image models. In USENIX Security.
@@
- Shao et al. (2023) Wenqi Shao, Yutao Hu, Peng Gao, Meng Lei, Kaipeng Zhang, Fanqing Meng, Peng Xu, Siyuan Huang, Hongsheng Li, Yu Qiao, et al. 2023. Tiny lvlm-ehub: Early multimodal experiments with bard. arXiv preprint arXiv:2308.03729 (2023).
@@
- Sharif et al. (2016) Mahmood Sharif, Sruti Bhagavatula, Lujo Bauer, and Michael K Reiter. 2016. Accessorize to a crime: Real and stealthy attacks on state-of-the-art face recognition. In CCS.
@@
- Sharif et al. (2019) Mahmood Sharif, Sruti Bhagavatula, Lujo Bauer, and Michael K Reiter. 2019. A general framework for adversarial examples with objectives. TOPS (2019).
@@
- Sharma et al. (2019) Yash Sharma, Gavin Weiguang Ding, and Marcus Brubaker. 2019. On the effectiveness of low frequency perturbations. arXiv preprint arXiv:1903.00073 (2019).
@@
- Sharma et al. (2018) Yash Sharma, Tien-Dung Le, and Moustafa Alzantot. 2018. Caad 2018: Generating transferable adversarial examples. arXiv preprint arXiv:1810.01268 (2018).
@@
- Shayegani et al. (2023) Erfan Shayegani, Yue Dong, and Nael Abu-Ghazaleh. 2023. Jailbreak in pieces: Compositional adversarial attacks on multi-modal language models. In ICLR.
@@
- Shen et al. (2023) Xinyue Shen, Zeyuan Chen, Michael Backes, Yun Shen, and Yang Zhang. 2023. "do anything now": Characterizing and evaluating in-the-wild jailbreak prompts on large language models. arXiv preprint arXiv:2308.03825 (2023).
@@
- Sid (2023) Sid. 2023. ChatGPT gives you free Windows 10 Pro keys! And it surprisingly works. (2023). https://x.com/immasiddtweets/status/1669721470006857729
@@
- Simonyan (2013) Karen Simonyan. 2013. Deep inside convolutional networks: Visualising image classification models and saliency maps. arXiv preprint arXiv:1312.6034 (2013).
@@
- Simonyan and Zisserman (2014) Karen Simonyan and Andrew Zisserman. 2014. Very deep convolutional networks for large-scale image recognition. arXiv preprint arXiv:1409.1556 (2014).
@@
- Song et al. (2018) Dawn Song, Kevin Eykholt, Ivan Evtimov, Earlence Fernandes, Bo Li, Amir Rahmati, Florian Tramer, Atul Prakash, and Tadayoshi Kohno. 2018. Physical adversarial examples for object detectors. In WOOT.
@@
- Stallkamp et al. (2012) Johannes Stallkamp, Marc Schlipsing, Jan Salmen, and Christian Igel. 2012. Man vs. computer: Benchmarking machine learning algorithms for traffic sign recognition. Neural networks (2012).
@@
- Su et al. (2019) Jiawei Su, Danilo Vasconcellos Vargas, and Kouichi Sakurai. 2019. One pixel attack for fooling deep neural networks. TEC (2019).
@@
- Su et al. (2023) Yixuan Su, Tian Lan, Huayang Li, Jialu Xu, Yan Wang, and Deng Cai. 2023. Pandagpt: One model to instruction-follow them all. arXiv preprint arXiv:2305.16355 (2023).
@@
- Sun et al. (2023b) Hao Sun, Zhexin Zhang, Jiawen Deng, Jiale Cheng, and Minlie Huang. 2023b. Safety assessment of chinese large language models. arXiv preprint arXiv:2304.10436 (2023).
@@
- Sun et al. (2023a) Quan Sun, Yuxin Fang, Ledell Wu, Xinlong Wang, and Yue Cao. 2023a. Eva-clip: Improved training techniques for clip at scale. arXiv preprint arXiv:2303.15389 (2023).
@@
- Sun et al. (2024) Quan Sun, Jinsheng Wang, Qiying Yu, Yufeng Cui, Fan Zhang, Xiaosong Zhang, and Xinlong Wang. 2024. Eva-clip-18b: Scaling clip to 18 billion parameters. arXiv preprint arXiv:2402.04252 (2024).
- Szegedy (2013) C Szegedy. 2013. Intriguing properties of neural networks. arXiv preprint arXiv:1312.6199 (2013).
@@
- Tabacof et al. (2016) Pedro Tabacof, Julia Tavares, and Eduardo Valle. 2016. Adversarial images for variational autoencoders. arXiv preprint arXiv:1612.00155 (2016).
@@
- Tan et al. (2024) Zhen Tan, Chengshuai Zhao, Raha Moraffah, Yifan Li, Yu Kong, Tianlong Chen, and Huan Liu. 2024. The Wolf Within: Covert Injection of Malice into MLLM Societies via an MLLM Operative. arXiv preprint arXiv:2402.14859 (2024).
- Thys et al. (2019) Simen Thys, Wiebe Van Ranst, and Toon Goedemé. 2019. Fooling automated surveillance cameras: adversarial patches to attack person detection. In CVPRW.
- Together (2023) Together. 2023. Releasing 3B and 7B RedPajama-INCITE family of models including base, instruction-tuned & chat models. (2023). https://www.together.ai/blog/redpajama-models-v1
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023a. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288 (2023).
- Tramèr et al. (2017a) Florian Tramèr, Alexey Kurakin, Nicolas Papernot, Ian Goodfellow, Dan Boneh, and Patrick McDaniel. 2017a. Ensemble adversarial training: Attacks and defenses. arXiv preprint arXiv:1705.07204 (2017).
- Tramèr et al. (2017b) Florian Tramèr, Nicolas Papernot, Ian Goodfellow, Dan Boneh, and Patrick McDaniel. 2017b. The space of transferable adversarial examples. arXiv preprint arXiv:1704.03453 (2017).
- Tseng et al. (2024) Kuo-Chun Tseng, Wei-Chieh Lai, Wei-Chun Huang, Yao-Chung Chang, and Sherali Zeadally. 2024. AI Threats: Adversarial Examples with a Quantum-Inspired Algorithm. IEEE Consumer Electronics Magazine (2024).
- Tsinghua (2023) Tsinghua. 2023. VisualGLM-6B. (2023). https://github.com/THUDM/VisualGLM-6B?tab=readme-ov-file
- Tu et al. (2023) Haoqin Tu, Chenhang Cui, Zijun Wang, Yiyang Zhou, Bingchen Zhao, Junlin Han, Wangchunshu Zhou, Huaxiu Yao, and Cihang Xie. 2023. How many unicorns are in this image? a safety evaluation benchmark for vision llms. arXiv preprint arXiv:2311.16101 (2023).
- Tunstall et al. (2023) Lewis Tunstall, Edward Beeching, Nathan Lambert, Nazneen Rajani, Kashif Rasul, Younes Belkada, Shengyi Huang, Leandro von Werra, Clémentine Fourrier, Nathan Habib, et al. 2023. Zephyr: Direct distillation of lm alignment. arXiv preprint arXiv:2310.16944 (2023).
- UnitaryAI (2020) UnitaryAI. 2020. Detoxify. (2020). https://github.com/unitaryai/detoxify
@@
- Walkerspider (2023) Walkerspider. 2023. DAN is my new friend. (2023). https://old.reddit.com/r/ChatGPT/comments/zlcyr9/dan_is_my_new_friend/
- Wang et al. (2024a) Haodi Wang, Kai Dong, Zhilei Zhu, Haotong Qin, Aishan Liu, Xiaolin Fang, Jiakai Wang, and Xianglong Liu. 2024a. Transferable multimodal attack on vision-language pre-training models. In S&P.
- Wang et al. (2020a) Haohan Wang, Xindi Wu, Zeyi Huang, and Eric P Xing. 2020a. High-frequency component helps explain the generalization of convolutional neural networks. In CVPR.
- Wang et al. (2024e) Ruofan Wang, Xingjun Ma, Hanxu Zhou, Chuanjun Ji, Guangnan Ye, and Yu-Gang Jiang. 2024e. White-box Multimodal Jailbreaks Against Large Vision-Language Models. arXiv preprint arXiv:2405.17894 (2024).
- Wang et al. (2024d) Siyuan Wang, Zhuohan Long, Zhihao Fan, and Zhongyu Wei. 2024d. From LLMs to MLLMs: Exploring the Landscape of Multimodal Jailbreaking. arXiv preprint arXiv:2406.14859 (2024).
- Wang et al. (2023a) Wenhui Wang, Hangbo Bao, Li Dong, Johan Bjorck, Zhiliang Peng, Qiang Liu, Kriti Aggarwal, Owais Khan Mohammed, Saksham Singhal, Subhojit Som, et al. 2023a. Image as a foreign language: Beit pretraining for vision and vision-language tasks. In CVPR.
- Wang et al. (2023c) Weihan Wang, Qingsong Lv, Wenmeng Yu, Wenyi Hong, Ji Qi, Yan Wang, Junhui Ji, Zhuoyi Yang, Lei Zhao, Xixuan Song, et al. 2023c. Cogvlm: Visual expert for pretrained language models. arXiv preprint arXiv:2311.03079 (2023).

### v2

@@ 9. Acknowledgments
- Niu et al. (2024) Zhenxing Niu, Haodong Ren, et al. 2024. Jailbreaking attack against multimodal large language model. arXiv:2402.02309.
- NousResearch (2023) NousResearch. 2023. Nous Hermes 2 - Yi-34B. https://huggingface.co/NousResearch/Nous-Hermes-2-Yi-34B
- OpenAI (2022) OpenAI. 2022. Moderation API. https://platform.openai.com/docs/guides/moderation/overview
- OpenAI (2023a) OpenAI. 2023a. GPT-4o API. https://platform.openai.com/docs/models/gpt-4o
- OpenAI (2023b) OpenAI. 2023b. Gpt-4v(ision) system card. https://cdn.openai.com/papers/GPTV_System_Card.pdf
- OpenAI (2023c) OpenAI. 2023c. Using GPT-4 for content moderation. https://openai.com/index/using-gpt-4-for-content-moderation/
- OpenAI (2024) OpenAI. 2024. OpenAI Usage policies. https://openai.com/policies/usage-policies
- OpenBMB (2024) OpenBMB. 2024. OmniLMM. https://github.com/OpenBMB/MiniCPM-V/blob/main/omnilmm_en.md
@@
- Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, et al. 2022. Training language models to follow instructions with human feedback. In NIPS.
@@
- Papernot et al. (2016a) Nicolas Papernot et al. 2016a. Towards the science of security and privacy in machine learning. arXiv:1611.03814.
- Papernot et al. (2016b) Nicolas Papernot, Patrick McDaniel, and Ian Goodfellow. 2016b. Transferability in machine learning: from phenomena to black-box attacks using adversarial samples. arXiv:1605.07277.
- Papernot et al. (2017) Nicolas Papernot, Patrick McDaniel, Ian Goodfellow, et al. 2017. Practical black-box attacks against machine learning. In ACM ASIACCS.
- Papernot et al. (2016c) Nicolas Papernot, Patrick McDaniel, Somesh Jha, et al. 2016c. The limitations of deep learning in adversarial settings. In EuroS&P.
@@
- Papineni et al. (2002) Kishore Papineni et al. 2002. Bleu: a method for automatic evaluation of machine translation. In ACL.
@@
- Peng et al. (2018) Yuxin Peng, Jinwei Qi, and Yuxin Yuan. 2018. Modality-specific cross-modal similarity measurement with recurrent attention network. TIP.
- Perez and Ribeiro (2022) Fábio Perez and Ian Ribeiro. 2022. Ignore previous prompt: Attack techniques for language models. arXiv:2211.09527.
- Phute et al. (2023) Mansi Phute, Alec Helbling, et al. 2023. Llm self defense: By self examination, llms know they are being tricked. arXiv:2308.07308.
- Pintor et al. (2021) Maura Pintor, Fabio Roli, Wieland Brendel, et al. 2021. Fast minimum-norm adversarial attacks through adaptive norm constraints. In NIPS.
- Plummer et al. (2015) Bryan A Plummer et al. 2015. Flickr30k entities: Collecting region-to-phrase correspondences for richer image-to-sentence models. In ICCV.
@@
- Qi et al. (2023) Xiangyu Qi et al. 2023. Fine-tuning aligned language models compromises safety, even when users do not intend to! arXiv:2310.03693.
- Qi et al. (2024) Xiangyu Qi, Kaixuan Huang, Ashwinee Panda, et al. 2024. Visual adversarial examples jailbreak aligned large language models. In AAAI.
- Qin et al. (2022) Zeyu Qin, Yanbo Fan, Yi Liu, et al. 2022. Boosting the transferability of adversarial attacks with reverse adversarial perturbation. In NIPS.
- Qraitem et al. (2024) Maan Qraitem, Nazia Tasnim, et al. 2024. Vision-llms can fool themselves with self-generated typographic attacks. arXiv:2402.00626.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, et al. 2021. Learning transferable visual models from natural language supervision. In ICML.
- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog.
- Rashtchian et al. (2010) Cyrus Rashtchian, Peter Young, et al. 2010. Collecting image annotations using amazon’s mechanical turk. In NAACL-HLT Workshop.
- Rasiwasia et al. (2010) Nikhil Rasiwasia, Jose Costa Pereira, Emanuele Coviello, et al. 2010. A new approach to cross-modal multimedia retrieval. In ACM MM.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, et al. 2022. High-resolution image synthesis with latent diffusion models. In CVPR.
- Rosen (1960) Jo Bo Rosen. 1960. The gradient projection method for nonlinear programming. Part I. Linear constraints. Journal of the society for industrial and applied mathematics.
- Röttger et al. (2023) Paul Röttger et al. 2023. Xstest: A test suite for identifying exaggerated safety behaviours in large language models. arXiv:2308.01263.
- Rudin et al. (1992) Leonid I Rudin, Stanley Osher, et al. 1992. Nonlinear total variation based noise removal algorithms. Physica D: nonlinear phenomena.
- Russakovsky et al. (2015) Olga Russakovsky, Jia Deng, Hao Su, Jonathan Krause, et al. 2015. Imagenet large scale visual recognition challenge. IJCV.
- Salman et al. (2019) Hadi Salman, Jerry Li, Ilya Razenshteyn, et al. 2019. Provably robust deep learning via adversarially trained smoothed classifiers. In NIPS.
@@
- Selvaraju et al. (2017) Ramprasaath R Selvaraju et al. 2017. Grad-cam: Visual explanations from deep networks via gradient-based localization. In CVPR.
- SessionGloomy (2023) SessionGloomy. 2023. New jailbreak! Proudly unveiling the tried and tested DAN 5.0. https://www.reddit.com/r/ChatGPT/comments/10tevu1/new_jailbreak_proudly_unveiling_the_tried_and/
- Shafahi et al. (2018) Ali Shafahi, W Ronny Huang, Mahyar Najibi, et al. 2018. Poison frogs! targeted clean-label poisoning attacks on neural networks. In NIPS.
- Shamir et al. (2021) Adi Shamir, Odelia Melamed, et al. 2021. The dimpled manifold model of adversarial examples in machine learning. arXiv:2106.10151.
- Shamsabadi et al. (2021) Ali Shahin Shamsabadi, Changjae Oh, and Andrea Cavallaro. 2021. Semantically adversarial learnable filters. TIP.
@@
- Shan et al. (2023) Shawn Shan, Jenna Cryan, et al. 2023. Glaze: Protecting artists from style mimicry by Text-to-Image models. In USENIX Security.
@@
- Shao et al. (2023) Wenqi Shao, Yutao Hu, Peng Gao, Meng Lei, et al. 2023. Tiny lvlm-ehub: Early multimodal experiments with bard. arXiv:2308.03729.
@@
- Sharif et al. (2016) Mahmood Sharif, Sruti Bhagavatula, et al. 2016. Accessorize to a crime: Real and stealthy attacks on state-of-the-art face recognition. In CCS.
@@
- Sharif et al. (2019) Mahmood Sharif, Sruti Bhagavatula, Lujo Bauer, et al. 2019. A general framework for adversarial examples with objectives. TOPS.
@@
- Sharma et al. (2019) Yash Sharma, Gavin Weiguang Ding, and Marcus Brubaker. 2019. On the effectiveness of low frequency perturbations. arXiv:1903.00073.
@@
- Sharma et al. (2018) Yash Sharma, Tien-Dung Le, and Moustafa Alzantot. 2018. Caad 2018: Generating transferable adversarial examples. arXiv:1810.01268.
@@
- Shayegani et al. (2023) Erfan Shayegani, Yue Dong, et al. 2023. Jailbreak in pieces: Compositional adversarial attacks on multi-modal language models. In ICLR.
@@
- Shen et al. (2024) Xinyue Shen et al. 2024. "do anything now": Characterizing and evaluating in-the-wild jailbreak prompts on large language models. CCS.
- Shokri et al. (2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov. 2017. Membership inference attacks against machine learning models. In SP.
@@
- Sid (2023) Sid. 2023. ChatGPT gives you free Windows 10 Pro keys! https://x.com/immasiddtweets/status/1669721470006857729
@@
- Simonyan (2013) Karen Simonyan. 2013. Deep inside convolutional networks: Visualising image classification models and saliency maps. arXiv:1312.6034.
@@
- Simonyan and Zisserman (2014) Karen Simonyan and Andrew Zisserman. 2014. Very deep convolutional networks for large-scale image recognition. arXiv:1409.1556.
@@
- Song et al. (2018) Dawn Song, Kevin Eykholt, Ivan Evtimov, et al. 2018. Physical adversarial examples for object detectors. In WOOT.
@@
- Stallkamp et al. (2012) Johannes Stallkamp et al. 2012. Man vs. computer: Benchmarking machine learning algorithms for traffic sign recognition. Neural Networks.
@@
- Su et al. (2019) Jiawei Su, Danilo Vasconcellos Vargas, and Kouichi Sakurai. 2019. One pixel attack for fooling deep neural networks. TEC.
@@
- Su et al. (2023) Yixuan Su, Tian Lan, Huayang Li, Jialu Xu, et al. 2023. Pandagpt: One model to instruction-follow them all. arXiv:2305.16355.
@@
- Sun et al. (2023b) Hao Sun, Zhexin Zhang, Jiawen Deng, et al. 2023b. Safety assessment of chinese large language models. arXiv:2304.10436.
@@
- Sun et al. (2023a) Quan Sun, Yuxin Fang, Ledell Wu, et al. 2023a. Eva-clip: Improved training techniques for clip at scale. arXiv:2303.15389.
@@
- Sun et al. (2024) Quan Sun, Jinsheng Wang, Qiying Yu, et al. 2024. Eva-clip-18b: Scaling clip to 18 billion parameters. arXiv:2402.04252.
- Szegedy et al. (2014) Christian Szegedy, Wojciech Zaremba, Ilya Sutskever, Joan Bruna, Dumitru Erhan, et al. 2014. Intriguing properties of neural networks. ICLR.
@@
- Tabacof et al. (2016) Pedro Tabacof, Julia Tavares, and Eduardo Valle. 2016. Adversarial images for variational autoencoders. arXiv:1612.00155.
@@
- Tan et al. (2024) Zhen Tan et al. 2024. The Wolf Within: Covert Injection of Malice into MLLM Societies via an MLLM Operative. arXiv:2402.14859.
- Tanay and Griffin (2016) Thomas Tanay and Lewis Griffin. 2016. A boundary tilting persepective on the phenomenon of adversarial examples. arXiv:1608.07690.
- Thys et al. (2019) Simen Thys, Wiebe Van Ranst, et al. 2019. Fooling automated surveillance cameras: adversarial patches to attack person detection. In CVPRW.
- Together (2023) Together. 2023. Releasing 3B and 7B RedPajama-INCITE family of models including base, instruction-tuned & chat models. https://www.together.ai/blog/redpajama-models-v1
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, et al. 2023a. Llama: Open and efficient foundation language models. arXiv:2302.13971.
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv:2307.09288.
- Tramèr et al. (2017a) Florian Tramèr, Alexey Kurakin, et al. 2017a. Ensemble adversarial training: Attacks and defenses. arXiv:1705.07204.
- Tramèr et al. (2017b) Florian Tramèr, Nicolas Papernot, et al. 2017b. The space of transferable adversarial examples. arXiv:1704.03453.
- Tramèr et al. (2016) Florian Tramèr, Fan Zhang, Ari Juels, Michael K Reiter, et al. 2016. Stealing machine learning models via prediction APIs. In USENIX Security.
- Tseng et al. (2024) Kuo-Chun Tseng et al. 2024. AI Threats: Adversarial Examples with a Quantum-Inspired Algorithm. IEEE Consumer Electronics Magazine.
- Tsinghua (2023) Tsinghua. 2023. VisualGLM-6B. https://github.com/THUDM/VisualGLM-6B?tab=readme-ov-file
- Tu et al. (2023) Haoqin Tu et al. 2023. How many unicorns are in this image? a safety evaluation benchmark for vision llms. arXiv:2311.16101.
- Tunstall et al. (2023) Lewis Tunstall, Edward Beeching, Nathan Lambert, et al. 2023. Zephyr: Direct distillation of lm alignment. arXiv:2310.16944.
- UnitaryAI (2020) UnitaryAI. 2020. Detoxify. https://github.com/unitaryai/detoxify
- Vadillo et al. (2025) Jon Vadillo, Roberto Santana, and Jose A Lozano. 2025. Adversarial attacks in explainable machine learning: A survey of threats against models and humans. Wiley Interdisciplinary Reviews: Data Mining and Knowledge Discovery.
@@
- Walkerspider (2023) Walkerspider. 2023. DAN is my new friend. https://old.reddit.com/r/ChatGPT/comments/zlcyr9/dan_is_my_new_friend/
- Wang et al. (2024a) Haodi Wang, Kai Dong, Zhilei Zhu, Haotong Qin, et al. 2024a. Transferable multimodal attack on vision-language pre-training models. In S&P.
- Wang et al. (2020a) Haohan Wang, Xindi Wu, et al. 2020a. High-frequency component helps explain the generalization of convolutional neural networks. In CVPR.
- Wang et al. (2024e) Ruofan Wang, Xingjun Ma, et al. 2024e. White-box Multimodal Jailbreaks Against Large Vision-Language Models. arXiv:2405.17894.
- Wang et al. (2024b) Siyuan Wang et al. 2024b. From LLMs to MLLMs: Exploring the Landscape of Multimodal Jailbreaking. arXiv:2406.14859.
- Wang et al. (2023a) Wenhui Wang, Hangbo Bao, Li Dong, et al. 2023a. Image as a foreign language: Beit pretraining for vision and vision-language tasks. In CVPR.
- Wang et al. (2023c) Weihan Wang, Qingsong Lv, Wenmeng Yu, et al. 2023c. Cogvlm: Visual expert for pretrained language models. arXiv:2311.03079.
- Wang et al. (2023b) Xunguang Wang et al. 2023b. InstructTA: Instruction-Tuned Targeted Attack for Large Vision-Language Models. arXiv:2312.01886.